OpenStudy (anonymous):

The rectangular floor of a storage building had a length of 25 yards and a perimeter of 90 yards. A construction crew knocked out one wall and extended the length of the floor by 10 yards. What was the area of the new floor? A. 1400 yd2 B. 700 yd2 C. 750 yd2 D. 500 yd2

OpenStudy (anonymous):

help me please

OpenStudy (calculusxy):

One length of the rectangle is = 25 Rectangles have two lengths and two widths. If we multiply 25*2=50 Our total length is 50yds. To make this problem easier to solve, I would suggest for you to find the width. A rectangle has two widths (2x). Our equation will be 2x+50=90. 90-50=40 2x=40 x=20 Since the length has been expanded by 10 yards we do (25+35). Our width we know is 20. Area=l*w and our length is 35 and our width is 20. So 35*20=700yd2.

OpenStudy (anonymous):

thx

OpenStudy (calculusxy):

Welcome